OpenStudy (anonymous):

PLEASE HELP!! Tyra is using broken pieces of tile to decorate the surface a small table. She is tiling the entire exterior surface of the table except for the part that touches the floor. The table consists of a cylindrical top with diameter fourteen inches and height 3 inches and a cubic base with side length 8 inches. How much area will Tyra have to tile? Use 3.14 for pi. Show your work.

OpenStudy (unklerhaukus):

can you draw it?

OpenStudy (anonymous):

That is what it looks like

OpenStudy (unklerhaukus):

i see; a circle and some rectangles

OpenStudy (anonymous):

Click the attachment

OpenStudy (anonymous):

cLick where it says 08_ac_07_013.jpg

OpenStudy (unklerhaukus):

the top surface is the 7in radius circle

OpenStudy (unklerhaukus):

there are four 8in x 8in rectangles

OpenStudy (unklerhaukus):

and one 3 x 2π7 rectangle

OpenStudy (anonymous):

show ur work plz

OpenStudy (anonymous):

for like the whole thing

OpenStudy (unklerhaukus):

if you also need to tile the underside of the table , that area will be the area of the 7in circle minus a 8in x 8in square

OpenStudy (anonymous):

SO whats the real anser?

OpenStudy (unklerhaukus):

you tell me

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

hi r u still on?

OpenStudy (unklerhaukus):

i'm still 'ere

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I hav a question

OpenStudy (anonymous):

What is tteh formula of the surface area of a cylinder?

OpenStudy (unklerhaukus):

well the surface of a cylinder is two circles plus a rectangle, \[SA_\text{cylinder}=2\times\pi r^2+2\pi r h\]

OpenStudy (anonymous):

Thx